# VisLTR: Visualization-in-the-Loop Table Reasoning

### v1

@@ 3 The Visualization-in-the-Loop Framework
Visualization Alignment. As the cornerstone of our framework, an LVLM is fine-tuned to establish a single joint embedding space for various modalities, including chart ($\mathbb{C}$), text ($\mathbb{T}$), and sketch ($\mathbb{S}$). To address the issue of a lack of suitable training data, we first curate a new dataset with chart-text pairings augmented from existing Chart-to-Text dataset[23], and chart-sketch pairings through user labeling. We apply pre-trained CLIP[41] to fine-tune a multi-modal model on our composite dataset. Moreover, we use a Transformer[42] for text modality, and ViT[10] encoders for chart and sketch modalities, respectively. To achieve alignment across modalities within the joint embedding space, we implement a bidirectional hinge-based triplet ranking loss function and a two-level cross-entropy loss function. Through this process, we successfully align the three modalities and yield strong performance across four evaluation metrics surpassing a threshold of 0.85.
@@
Visualization Referencing. In the visualization referencing module, we generate visualization references that comprehensively capture key data patterns inherent in an input table. We adopt a two-stage approach that first decomposes the input table $D$ into data facets, which are subsequently mapped to visualization references. In the facet generation stage, we employ a multifaceted process involving smoothing to preserve critical patterns while eliminating noise, as well as segmentation to extract meaningful and reliable patterns. Next in the visual mapping stage, these facets are mapped into fixed-size visualization references, effectively representing the patterns within the original tabular data.
@@
Visualization Pruning. In the visualization pruning module, we introduce pruning and indexing techniques to enable timely visualization reference retrieval. Specifically, visualization pruning works by filtering out less informative visualization references among those with a high degree of similarity, thereby reducing the number of stored visualization references meanwhile keeping key data change patterns. In parallel, visualization indexing employs an open-source vector database, Chroma[1], to efficiently store and retrieve visualization references. Here, we use the fine-tuned LVLM to transform all the pruned visualization references into 512-dimensional vectors, which are then indexed by Chroma for rapid and accurate follow-up retrieval. This combination of pruning and indexing ensures VisLTR can swiftly manage and retrieve massive visualization references with precision.
@@
Visualization Interaction. In the visualization interaction module, we introduce a “decomposing-execution-filling” strategy to enable the connection of multi-modal user queries with the stored embeddings of visualization references. Specifically, the user query is first decomposed into a visualization retrieval task. Second, the retrieval task is executed in Chrome to get a retrieved visualization reference. Afterward, with the help of the fine-tuned LVLM, the corresponding trend word of the retrieved visualization reference is generated and then filled into the textual response to the user. An interactive visual interface with multiple views and rich interactions is also developed to support the iterative table reasoning process.
4 Multi-modal Visualization Alignment
@@
We set out to establish an LVLM that aligns visualization modalities of chart, sketch, and text in a single joint embedding space. This section presents the data preparation, including data augmentation and user labeling strategies, to address the issue of lacking training dataset (Sect. 4.1), followed by modality binding for establishing a joint embedding space that binds chart, sketch, and text modalities together (Sect. 4.2). In the end, we present the implementation details (Sect. 4.3).
@@ 4.1 Data Preparation
To train the multi-modal visualization alignment model, we need to first prepare a dataset containing user intentions, which is not covered by current open-sourced datasets. To achieve the alignment of sketches, charts, and text, we employ two strategies to construct the training data: (1) data augmentation for chart-text pairing, and (2) user labeling for chart-sketch pairing. Figure 2 uses an example to illustrate the two data preparation strategies.
@@ 4.2 Modality Binding
Given a chart $C_{i}$ with its corresponding observation $T_{i}$ in text modality, and observation $S_{i}$ in sketch modality, we encode them into normalized embeddings: $p_{i}=f(C_{i})$, $q_{i}=f(S_{i})$, and $k_{i}=f(T_{i})$, where $f$ is a 63M-parameter 12-layer 512-wide Transformer with the architecture modifications with 8 attention heads[42] for text modality, and the Vision Transformer (ViT)[10] with a patch size of 32 for chart and sketch modalities. Each modality is processed by a separate encoder, which is designed to cater to the unique characteristics of the input. We add a modality-specific linear projection head on each encoder to obtain a fixed size $d$ dimensional embedding, which is normalized, and set as $p_{i}$, $q_{i}$ and $k_{i}$ for loss calculation. Besides facilitating learning, this setup allows us to initialize a subset of the encoders with weights from pre-trained CLIP[41] models. This initialization not only provides a strong starting point for our LVLM but also enables it to better understand and align the semantic content across different modalities.
@@
where $\alpha$ denotes the margin parameter, and $[x]_{+}\equiv\max(x,0)$. $C^{\prime}=\arg\max_{X\neq C}H_{2}(X,S)$ and $S^{\prime}=\arg\max_{Y\neq S}H_{2}(C,Y)$ denote the hardest negative corresponding to the positive pair $(C,S)$. The loss makes the chart embedding $p_{i}$ and the sketch embedding $q_{i}$ closer in the joint embedding space, and thus aligns $\mathbb{C}$ and $\mathbb{S}$. Set $L$ as the final loss of three modalities in the embedding space, that is, the sum of the two ranking losses:
$L(T,C,S)=L_{1}(C,T)+L_{2}(C,S).$
By minimizing both the loss between chart and text modalities, and the loss between chart and sketch modalities, we achieve the alignment between the similar embeddings of the three modalities in a joint embedding space.
@@ 4.3 Implementation Details
5 VisLTR System
With the LVLM for multi-modal visualization alignment, we build VisLTR system that allows users to intuitively explore, comprehend, and interact with tabular data using visualizations as a proxy. This section details the implementation of visualization referencing (Sect. 5.1), visualization pruning (Sect. 5.2), and visualization interaction (Sect. 5.3) in VisLTR system.
@@ 5.1 Visualization Referencing
Facet Generation. In the first stage, we aim to capture meaningful data facets of the input table as comprehensively as possible. If we use violent enumeration, it will produce a lot of candidate data facets, and many of them will not include data trends. To avoid the inefficient data facet generation, we adopt the idea of subset embedding [54] indicating that patterns of multi-dimensional data are often related to subsets rather than records. Specifically, we adopt a multifaced process that includes a dual strategy of smoothing and segmentation.
The multifaceted process we adopt incorporates a dual strategy focused on smoothing and segmentation. We use a time-series table as an instance for explanation, as it is rich in data change patterns that SQL-based LLM for table reasoning is not able to address. The initial step involves smoothing different variables to retain pivotal patterns while filtering out the noise. To achieve this, VisLTR employs Gaussian Smoothing, a technique favored for its property of maintaining shape consistency under the Fourier transform. Compared with widely used Moving Average (MA) or Weighted Moving Average (WMA)[18], Gaussian Smoothing can not only preserve key features while smoothing but also have flexibility in parameter control.
After smoothing, we focus on segmenting the table $D(d_{1},d_{2},\cdots,d_{n})$ into facets to extract meaningful and reliable patterns. In order to facilitate the discussion, we focus on the segmentation of one variable $d_{i}$ in the following. This is accomplished by identifying a set of data facets $r(d_{i})$ for each variable $d_{i}$, which represents a continuous subset of values. To ensure each data facet reflects a valid data pattern, we incorporate the Page-Hinckley Test (PHT)[38] to identify critical points that demote substantial changes. By setting these critical points as boundaries for the segments, we can delineate different facets of the variable $d_{i}$. The comprehensive set of data facets for the variable $d_{i}$ is represented by $r(d_{i})=PHT(d_{i})$.
Visual Mapping. The second stage is visual mapping, where we harness the matplotlib library to transform each data facet into a fixed-size visualization reference that is devoid of axes, specifically utilizing bar, line, and area charts to manifest the trend in each data facet. Furthermore, the uniformity provided by fixed-size visualization references improves the fine-tuned LVLM’s capacity to analyze and process them. The visualization references are represented as $C=(c_{1},\cdots,c_{n})$, where $c_{j}=vmap(r_{j}(d_{i}))$, where $r_{j}(d_{i})$ is a data facet selected from the set $r(d_{i})$, and $vmap$ represents the visual mapping process.
@@ 5.2 Visualization Pruning and Indexing
For a given table, we will generate a vast collection of visualization references, each of which is subsequently mapped into vector representations (i.e., visualization embeddings). The large number of visualization references poses significant challenges to efficient storage and retrieval. To address these challenges, we first introduce a visualization pruning technique to effectively reduce the overall number of visualization references, thereby reducing the required storage space and initially streamlining the retrieval process. However, even after pruning, the remaining number of visualization references may still be too large, demanding a visualization indexing technique to accelerate the retrieval process. The combined approach of visualization pruning and indexing, as illustrated in Figure 4, significantly improves VisLTR’s ability to manage and retrieve visualization references.
Visualization Pruning. The key idea of visualization pruning is to filter out less informative visualization references in advance, narrowing down the pool of candidates to those most likely to match the query. The main challenge is retaining highly valuable visualization references during pruning. To this end, we introduce a visualization pruning method to filter out less informative visualization references. For all visualization references for a variable $d_{i}$, it may encounter a scenario where two references exhibit similar change patterns, and one is entirely encompassed by the other. An example scenario is illustrated in Figure 4, where the ‘a peak’ change pattern exists in pink-boxed visualization references. In this case, the light-pink-boxed one has less information and is consequently filtered out. By leveraging the visualization pruning technique, we filtered out about two-thirds of the visualizations that contained less information, effectively reducing the storage footprint.
Visualization Indexing. For efficient retrieval of visualization references, we utilize an open-sourced vector database named Chroma[1], which enables effective indexing and retrieval of embeddings. Specifically, we first apply our fine-tuned LVLM to encode all visualization references into 512-dimensional vectors. Chroma automatically maintains an index document when encoding all visualization references for retrieval. These indexes compress vectors into codes of a fixed size, storing them in an array. For similarity measurement, we employ the dot product function. By leveraging Chroma’s advanced indexing capabilities, VisLTR efficiently speeds up the retrieval process to handle vast amounts of visualizations, ensuring both speed and accuracy.
@@
Figure 4: The pipeline of visualization referencing and visualization pruning modules. Facing the massive generated visualization references, two techniques including, visualization pruning and indexing are used to accelerate the storage and retrieval process.
@@
Figure 5: The interactive interface of VisLTR, which contains four main views and supports multi-modal interactions for table reasoning.
@@ 5.3 Visualization Interaction
The Decomposing-Execution-Filling Strategy. Inspired by Dater[56], we introduce a ‘decomposing-execution-filling’ strategy to decompose and parse multi-modal user queries and return results that meet user expectations. Specifically, we first decompose a user’s table reasoning queries and convert them into visualization reference retrieval operations, similar to text-to-SQL parsing (e.g.,[40, 48]). Afterward, we execute the visualization retrieval operations in Chroma and get the result for back-filling the final response.
@@
Here, we employ two LLMs with in-context learning using prompts to decompose user queries and complete outputs respectively. Below we illustrate an example using the case shown in Figure VisLTR: Visualization-in-the-Loop Table Reasoning(A). Given the query “What is the price trend of Apple during March?”, we first decompose the question into a retrieval task “return the visualization reference of Apple price in March”, and a cross-modal alignment task “identify the trend in the retrieved visualization reference”. The corresponding visualization reference is first retrieved from Chroma, and then aligned with the ‘two peaks’ pattern to generate the final textual response “There is a two-peak price trend in Apple during March.”. Detailed prompts are provided in our open-sourced code.
@@
Visual Interface. In addition to textual responses, VisLTR also incorporates a user-friendly interface to enable intuitive table reasoning and exploration. As shown in Figure 5, the interface includes four main views: (A) Data Table, (B) Main View, (C) Chat Box, and (D) Pattern View. When a user uploads a table, the back-end of VisLTR processes the tabular data into visualization references and stores them in the vector database, while the front-end interface displays the uploaded table in Data Table, and presents the visualization for the data in the Main View. By default, Main View uses an area graph for the overall context information and a line chart for the details. Various operations, such as selecting different variables or combinations of variables to display, are enabled in the drop-down box in the upper right corner of Main View. For the display of multi-variables, we use a multi-line chart with colors to encode different variables. The bottom area chart also allows for fixed-scale zooming and brushing, allowing users to observe detailed variable changes.
@@
The Chat Box provides a dialogue window that supports multi-round conversations and multi-modal inputs including uploading charts, drawing sketches, and typing text. Textual responses to user queries are also displayed in the Chat Box. The responses are also connected to Main View, where the corresponding visualization reference is highlighted. Pattern View presents groups of typical data change patterns for the variable selected in Main View. These groups are ordered from top to bottom, with the most populated data change pattern at the top, taping to the least at the bottom. Users can select a specific group of data change patterns, and then the top-3 visualization references of this pattern will be overlayed by highlighting the background. If there is an overlap among the displayed visualization references, the overlapping area corresponds to a darker background color. For users wishing to delve into each reference in detail, users can click the ‘show’ button located in the upper right corner of the Main View to achieve a circular review of each one. Additionally, Pattern View provides multi-modal information (description text, different chart types, and candidate sketches) for a single variable’s typical change patterns, including descriptive text, various charts, and potential sketch candidates.

### v2

@@ 3 The Visualization Enhanced Framework
Visualization Alignment. As the cornerstone of our framework, a MLLM is fine-tuned to establish a single joint embedding space for various modalities, including chart ($\mathbb{C}$), text ($\mathbb{T}$), and sketch ($\mathbb{S}$). To address the issue of a lack of suitable training data, we first curate a new dataset with chart-text pairings augmented from existing Chart-to-Text dataset[28], and chart-sketch pairings through user labeling. Utilizing the pre-trained CLIP[48], we fine-tune a multimodal LLM on our composite dataset, employing a Transformer[49] for text modality, and ViT[13] encoders for chart and sketch modalities, respectively. To achieve alignment across modalities within the joint embedding space, we implement a bidirectional hinge-based triplet ranking loss function and a two-level cross-entropy loss function. Through this process, we successfully align the three modalities and yield strong performance across four evaluation metrics surpassing a threshold of 0.85. A detailed description of the module is provided in Sect. 4.
@@
Visualization Referencing. In the visualization referencing module, we generate visualization references that comprehensively capture key data patterns inherent in an input table. This module employs a two-stage approach that first decomposes the input table $D$ into data facets, which are then mapped into fixed-size visualization references. The description of this module is detailed in Sect. 5.1.
@@
Visualization Pruning. In the visualization pruning module, we introduce pruning and indexing techniques to enable timely visualization reference retrieval. Specifically, visualization pruning works by filtering out less informative visualization references by measuring the similarity between visualization vectors using Euclidean distance and applying a pruning threshold. This process reduces the number of stored references while preserving essential data change patterns. Concurrently, visualization indexing leverages an open-source vector database, Chroma[1], to efficiently store and retrieve visualization references. Here, we use the fine-tuned MLLM to transform all the pruned visualization references into 512-dimensional vectors, which are then indexed by Chroma for rapid and accurate follow-up retrieval. This combination of pruning and indexing ensures VisTR can swiftly manage and retrieve massive visualization references with precision. This module is detailed in Sect. 5.2.
@@
Visualization Interaction. In the visualization interaction module, we introduce a “decomposing-execution-filling” strategy to enable the connection of multimodal user queries with the stored embeddings of visualization references. Specifically, the user query is first decomposed into a visualization retrieval task. Second, the retrieval task is executed in Chrome to get a retrieved visualization reference. Afterward, with the help of the fine-tuned MLLM, the corresponding trend word of the retrieved visualization reference is generated and then filled into the textual response to the user. An interactive visual interface with multiple views and rich interactions is also developed to support the iterative table reasoning process, transitioning from abstract to detailed exploration, as detailed in Sect. 5.3.
4 Multimodal Visualization Alignment
@@
We set out to establish a MLLM that aligns visualization modalities of chart, sketch, and text in a single joint embedding space. This section presents the data preparation, including data augmentation and user labeling strategies, to address the issue of lacking training dataset (Sect. 4.1), followed by modality binding for establishing a joint embedding space that binds chart, sketch, and text modalities together (Sect. 4.2). In the end, we present the implementation details (Sect. 4.3).
@@ 4.1 Data Preparation
To train the multimodal visualization alignment model, we need to first prepare a dataset containing user intentions, which is not covered by current open-sourced datasets. To achieve the alignment of sketches, charts, and text, we employ two strategies to construct the training data: (1) data augmentation for chart-text pairing, and (2) user labeling for chart-sketch pairing. Figure 2 uses an example to illustrate the two data preparation strategies.
@@ 4.2 Modality Binding
Given a chart $C_{i}$ with its corresponding observation $T_{i}$ in text modality, and observation $S_{i}$ in sketch modality, we encode them into normalized embeddings: $p_{i}=f(C_{i})$, $q_{i}=f(S_{i})$, and $k_{i}=f(T_{i})$, where $f$ is a 63M-parameter 12-layer 512-wide Transformer with the architecture modifications with 8 attention heads[49] for text modality, and the Vision Transformer (ViT)[13] with a patch size of 32 for chart and sketch modalities. Each modality is processed by a separate encoder, which is designed to cater to the unique characteristics of the input. We add a modality-specific linear projection head on each encoder to obtain a fixed size $d$ dimensional embedding, which is normalized, and set as $p_{i}$, $q_{i}$, and $k_{i}$ for loss calculation. Besides facilitating learning, this setup allows us to initialize a subset of the encoders with weights from pre-trained CLIP[48] models. This initialization not only provides a strong starting point for our MLLM but also enables it to better understand and align the semantic content across different modalities.
@@
where $\alpha$ denotes the margin parameter, and $[x]_{+}\equiv\max(x,0)$. $C^{\prime}=\arg\max_{X\neq C}H_{2}(X,S)$ and $S^{\prime}=\arg\max_{Y\neq S}H_{2}(C,Y)$ denote the hardest negative corresponding to the positive pair $(C,S)$. The loss makes the chart embedding $p_{i}$ and the sketch embedding $q_{i}$ closer in the joint embedding space, and thus aligns $\mathbb{C}$ and $\mathbb{S}$. Set $L$ as the final loss of three modalities in the embedding space, that is, the sum of the two ranking losses: $L(T,C,S)=L_{1}(C,T)+L_{2}(C,S).$ By minimizing both the loss between chart and text modalities, and the loss between chart and sketch modalities, we achieve the alignment between the similar embeddings of the three modalities in a joint embedding space.
@@ 4.3 Implementation Details
5 VisTR System
With the MLLM for multimodal visualization alignment, we build VisTR system that allows users to intuitively explore, comprehend, and interact with tabular data using visualizations as the representation. This section details the implementation of visualization referencing (Sect. 5.1), visualization pruning (Sect. 5.2), and visualization interaction (Sect. 5.3) in VisTR system.
@@ 5.1 Visualization Referencing
Facet Generation. In the first stage, we aim to capture meaningful data facets of the input table as comprehensively as possible. If we use violent enumeration, it will produce a lot of candidate data facets, and many of them will not include data trends. To avoid the inefficient data facet generation, we adopt the idea of subset embedding [62] indicating that patterns of multi-dimensional data are often related to subsets rather than records. Specifically, we adopt a multifaced process that incorporates a dual strategy of smoothing and segmentation.
Consider an input time-series table $D(timestamp,d_{1},d_{2},\cdots,d_{n})$, where each $d_{i}$ represents a distinct variable containing a sequence of data points indexed by the timestamp. The first step is to smooth each variable $d_{i}$ to retain essential patterns while filtering out the noise. To achieve this, VisTR employs Gaussian Smoothing, a technique favored for maintaining shape consistency under the Fourier transform[53]. Formally, for each each variable $d_{i}$, the smoothed series $\widetilde{d_{i}}$ is obtained as: $\widetilde{d_{i}}=GaussianSmoothing(d_{i})$. Compared with widely used Moving Average (MA) or Weighted Moving Average (WMA)[23], Gaussian Smoothing can not only preserve key features while smoothing but also have flexibility in parameter control.
After smoothing, we segment each smoothed variable $\widetilde{d_{i}}$ into distinct facets to extract meaningful and reliable patterns. The segmentation process is defined as $R(\widetilde{d_{i}})=\{r_{1}(\widetilde{d_{i}}),r_{2}(\widetilde{d_{i}}),\cdots r_{m}(\widetilde{d_{i}})\}$, where $r_{j}(\widetilde{d_{i}})$ represents a subset (termed as a facet) of $\widetilde{d_{i}}$. To ensure each data facet reflects a valid data pattern, we incorporate the Page-Hinckley Test (PHT)[44] to detect critical points where shifts occur, thereby delineating the boundaries between different facets. Specifically, the segmentation process can be formalized as $R(\widetilde{d_{i}})=PHT(\widetilde{d_{i}})$. Additionally, we aggregate continuous subsets to form new facets, enabling the inclusion of segments with varying lengths.
Visual Mapping. The second stage is visual mapping, where we transform each data facet into a standard visual representation. This process leverages the matplotlib library to generate fixed-size visualization references devoid of axes, specifically utilizing bar, line, and area charts to manifest the trend within each data facet. The uniformity provided by fixed-size visualization references improves the fine-tuned MLLM’s capacity to analyze and process them. For each data facet $r_{j}(\widetilde{d_{i}})\in R(\widetilde{d_{i}})$, the visual mapping process transforms $r_{j}(\widetilde{d_{i}})$ into a fixed-size visualization reference $c_{j}$, which is formalized as $c_{j}=vmap(r_{j}(\widetilde{d_{i}}))$. The collection of all visualization references derived from the data facets of a particular variable $d_{i}$ is represented as $C(d_{i})=\{c_{1},c_{2},\cdots,c_{m}\}$. Collectively, for all variables in the input table $D$, the visualization references are denoted as $C=\bigcup_{i=1}^{n}C(d_{i})=\{c_{i,j}|d_{i}\in D,r_{j}(d_{i})\in R(d_{i})\}$.
@@ 5.2 Visualization Pruning and Indexing
For a given table, we will generate a vast collection of visualization references, each of which is subsequently mapped into vector representations (i.e., embeddings). The large number of visualization references poses significant challenges to efficient storage and retrieval. To address these challenges, we design a combined approach of visualization pruning and indexing, as illustrated in Figure 4, to improve VisTR’s ability to manage and retrieve visualization references.
@@
Figure 4: The pipeline of visualization referencing and visualization pruning modules. Visualization pruning and indexing are used to accelerate the storage and retrieval process.
Visualization Pruning. Here, the key idea is to filter out less informative visualization references in advance while retaining those that provide significant insights. The main challenge lies in discerning and preserving highly valuable visualization references during pruning. For all visualization references for a variable $d_{i}$, it may encounter a scenario where two references exhibit similar change patterns, and one is entirely encompassed by the other. We quantify the similarity between visualization vectors using the Euclidean distance $d(c_{j},c_{k})$ between the two vectors $c_{j}$ and $c_{k}$. We introduce a pruning threshold to determine whether two visualization vectors are sufficiently similar to warrant pruning. If $d(c_{j},c_{k})$ is less than the set pruning threshold, the vector covering a shorter time range is filtered out. An example scenario is illustrated in Figure 4, where the ‘a peak’ change pattern exists in pink-boxed visualization references. In this case, the light-pink-boxed one has less information and is consequently filtered out.
Figure 5: The relationship between the number of visualization references, table rows, and the pruning threshold. The remaining visualization references are close to approximately 20% of the raw number when the pruning threshold is set to 1.
The pruning threshold acts as a predefined parameter that affects the number of stored visualization references. To assess its impact, we designed an experiment with four values: 0.5, 0.8, 1, and 1.5, and applied the thresholds to tables in a range of row numbers. Specifically, the row numbers were considered based on the typical sizes used in existing LLM-based table reasoning methods, which typically support data tables with fewer than 50 rows. For instance, Tabfact[7], a commonly used dataset containing over 16k Wiki tables, has an average of 12.96 rows per table (max.=48, min.=1, std.=8.46). Given the focus on time-series data tables and long-term trend analysis, we extended the experiment to tables with up to 1000 rows, approximately encompassing four trading years. We calculated the average number of visualization references across three sub-tables with corresponding table sizes extracted from the DJI. data table for each pruning threshold.
Figure 5 shows the relationship between the number of stored visualization references, the number of data table rows, and the pruning thresholds. When the pruning threshold is set to 1, the number of stored visualization references approximated 20% of the raw reference number. This aligns with the well-known Pareto principle (80/20 rule)[45], which is widely leveraged in various optimization strategies to balance efficiency and information retention. As such, we also set the pruning threshold to be 1. This strikes a balance between maintaining a rich representation set and reducing the storage footprint.
Visualization Indexing. For efficient retrieval of visualization references, we utilize an open-sourced vector database named Chroma[1], which enables effective indexing and retrieval of embeddings. Specifically, we first apply our fine-tuned MLLM to encode all visualization references into 512-dimensional vectors. Chroma automatically maintains an index document when encoding all visualization references for retrieval. These indexes compress vectors into codes of a fixed size, storing them in an array. For similarity measurement, we employ the dot product function. By leveraging Chroma’s advanced indexing capabilities, VisTR efficiently speeds up the retrieval process to handle vast amounts of visualizations, ensuring both speed and accuracy.
@@
Figure 6: The interactive interface of VisTR, which contains four main views and supports multimodal interactions for table reasoning.
@@ 5.3 Visualization Interaction
The Decomposing-Execution-Filling Strategy. Inspired by Dater[65], we introduce a ‘decomposing-execution-filling’ strategy to decompose and parse multimodal user queries and return results that meet user expectations. Specifically, we first decompose a user’s table reasoning queries and convert them into visualization reference retrieval operations, similar to text-to-SQL parsing (e.g.,[47, 56]). Afterward, we execute the visualization retrieval operations in Chroma and get the result for back-filling the final response.
@@
Here, we employ two LLMs with in-context learning using prompts to decompose user queries and complete outputs respectively. Below we illustrate an example using the case shown in Figure VisTR: Visualizations as Representations for Time-series Table Reasoning(A). Given the query “What is the price trend of Apple during March?”, we first decompose the question into a retrieval task “return the visualization reference of Apple price in March”, and a cross-modal alignment task “recognize the trend in the retrieved visualization reference”. The corresponding visualization reference is first retrieved from Chroma, and then aligned with the ‘two peaks’ pattern to generate the final textual response “There is a two-peak price trend in Apple during March.”. Detailed prompts are provided in our open-sourced code.
@@
Visual Interface. In addition to textual responses, VisTR also incorporates a user-friendly interface to enable intuitive table reasoning and exploration. As shown in Figure 6, the interface includes four main views: (A) Data Table, (B) Main View, (C) Chat Box, and (D) Pattern View. When a user uploads a table, the back-end of VisTR processes the tabular data into visualization references and stores them in the vector database, while the front-end interface displays the uploaded table in Data Table, and presents the visualization for the data in the Main View. By default, Main View uses an area graph for the overall context information and a line chart for the details. Various operations, such as selecting different variables or combinations of variables to display, are enabled in the drop-down box in the upper right corner of Main View. For the display of multi-variables, we use a multi-line chart with colors to encode different variables. The bottom area chart also allows for fixed-scale zooming and brushing, allowing users to observe detailed variable changes.
@@
The Chat Box provides a dialogue window that supports multi-round conversations and multimodal inputs including uploading charts, drawing sketches, and typing text. Textual responses to user queries are also displayed in the Chat Box. The responses are also connected to Main View, where the corresponding visualization reference is highlighted. Pattern View presents groups of typical data change patterns for the variable selected in Main View. These groups are ordered from top to bottom, with the most populated data change pattern at the top, taping to the least at the bottom. Users can select a specific group of data change patterns, and then the top-3 visualization references of this pattern will be overlayed by highlighting the background. If there is an overlap among the displayed visualization references, the overlapping area corresponds to a darker background color. For users wishing to delve into each reference in detail, users can click the ‘show’ button located in the upper right corner of the Main View to achieve a circular review of each one. Additionally, Pattern View provides multimodal information (description text, different chart types, and candidate sketches) for a single variable’s typical change patterns, including descriptive text, various charts, and potential sketch candidates.